OpenStudy (anonymous):

evaluate the limit 1. as x approaches 0 , (3x^2-x-2)/(7x^2+5x-1) 2. as x approaches 0 (tanx-sinx)/x^3

OpenStudy (anonymous):

for the first one i got 2 but i'm not sure of the answer

OpenStudy (freckles):

well you are write since (3(0)^2-0-2)=-2 and (7(0)^2)+5(0)-1)=-1 and -2/-1 is 2

OpenStudy (freckles):

right *

OpenStudy (anonymous):

okay , i got the first right.

OpenStudy (freckles):

The next one isn't as simple because you cannot just plug in 0.

OpenStudy (freckles):

because you get 0/0

OpenStudy (freckles):

0/0 means you have more work to do

OpenStudy (freckles):

do you know l'hopital's rule?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

that is differentiating the numerator and the denominator

OpenStudy (freckles):

yep

OpenStudy (anonymous):

wiii it be\[\frac{ \sec ^{2}x-cosx }{ 3x ^{2} }\]

OpenStudy (freckles):

yes but what happens when you plug in 0?

OpenStudy (anonymous):

it will still give 0/0

OpenStudy (freckles):

so do l'hoptal again

OpenStudy (freckles):

if we keep differentiating a polynomial we will eventually get a constant so don't worry this has a send

OpenStudy (freckles):

this will end*

OpenStudy (freckles):

this has a end it what I meant

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

so i have to keep on differentiating until i get a constant

OpenStudy (freckles):

well you can keep on using l'hoptal if you have the 0/0 eventually you won't have that because eventually the bottom will be equal to a constant that isn't 0